OpenStudy (anonymous):

The greater of two number is one less than 8 times the smaller. Their sum is 98. Find the numbers and equation.

OpenStudy (campbell_st):

let the smaller number be x the larger number is 8x - 1 so x + 8x - 1 = 98 now solve for x, you will get the smaller number and should be able to then calculate the larger number

OpenStudy (anonymous):

@campbell_st x=9

OpenStudy (campbell_st):

nope.... try again

OpenStudy (campbell_st):

if you simplify the equation you get 9x - 1 = 98 check your solution

OpenStudy (anonymous):

@campbell_st sorry i meant to say 11

OpenStudy (campbell_st):

thats correct.... now you can find the larger number

OpenStudy (anonymous):

8x-1=87

OpenStudy (campbell_st):

thats correct... well done

OpenStudy (anonymous):

@campbell_st No thank you :)